# Efficient Quantum Pseudorandomness from Hamiltonian Phase States

### v1

@@ 6.6.1 Security of the $\mathsf{FHFHFC}$ pseudo-random unitary
$\sum_{w,y\in\{0,1\}^{n}}(-1)^{\braket{x\oplus w,y}}\ket{w}\mathbb{E}_{\Theta_{1},\Theta_{2},\Theta_{3}}\alpha_{R}^{\Theta_{1},\Theta_{2},\Theta_{3}}\\ \exp\left(i\sum_{z}\theta_{z}^{(1)}(-1)^{\braket{x,z}}+\theta_{z}^{(2)}(-1)^{\braket{y,z}}+\theta_{z}^{(3)}(-1)^{\braket{w,z}}\right)\ket{\Theta_{1},\Theta_{2},\Theta_{3}}\,.$
Now, summing over $y$ (since it does not appear in the ket), we enforce the condition that $w=y\oplus x$ (otherwise the coefficients add to 00). Thus, this state is equal to the following.
$\sum_{y\in\{0,1\}^{n}}\ket{y}\mathbb{E}_{\Theta_{1},\Theta_{2},\Theta_{3}}\alpha_{R}^{\Theta_{1},\Theta_{2},\Theta_{3}}\exp\left(i\sum_{z}\theta_{z}^{(1)}(-1)^{\braket{x,z}}+\theta_{z}^{(2)}(-1)^{\braket{y,z}}+\theta_{z}^{(3)}(-1)^{\braket{x\oplus y,z}}\right)\ket{\Theta_{1},\Theta_{2},\Theta_{3}}\,.$
Expanding out the definition of $\alpha_{R}^{\Theta_{1},\Theta_{2},\Theta_{3}}$ and combining the product of exponentials yields the desired state. Since the $\mathsf{FHFHF}$ oracle is a unitary, the resulting state must be normalized, which yields the pre-factor of $\sqrt{2^{n}-|R|}$. ∎
We make one important note: While it may seem like one could apply a Hadamard to |$\ket{y}$ and recover the input $x$, |$\ket{y}$ is actually entangled to the purifying register, and applying a Hadamard will simply yield another random bit string, instead of un-doing the original Hadamard in the oracle.
Now, we have shown that the relation states for the $\mathsf{FHF}$ oracle are orthogonal, and that the action of the $\mathsf{FHF}$ oracle is to output a random $y$ and store $(x,y)$ in the relation state. Now, we make the connection to the path recording oracle formal by defining two isometries that will act on the purifying register, $\mathsf{Compress}$ and $\mathsf{Biject}$. The compress isometry acts as follows:
$\mathsf{Compress}=\sum_{R\ \mathrm{bijection}}\ket{R}\!\!\bra{\mathsf{fhfhf}_{R}}\,.$ | .
@@
$\mathsf{Biject}=\sum_{R\ \mathrm{bijection}}|\mathsf{fhfhf}_{R}\rangle\!\langle\mathsf{fhfhf}_{R}|\,.$
Let $\Pi^{(t)}$ be the projection onto bijection relation states, as defined in [ma2024pseudorandom]. Then we have the following lemma.
@@
$\mathsf{Compress}\cdot\mathsf{Biject}\ket{\mathcal{A}_{t}^{\mathsf{FHFHFO}\cdot G}}=\Pi^{(t)}\ket{\mathcal{A}^{\mathsf{PR}\cdot G}_{t}}\,.$
@@
Similar to the proof from [ma2024pseudorandom], we can explicitly write out the states |$\ket{\mathcal{A}_{t}^{\mathcal{O}}}$ in either case. For a set $S$, let $S^{t}$ denote the $t$-fold cartestian product of the set, and $S^{t}_{\mathrm{dist}}$ be the distinct subset of $S^{t}$ (i.e. all sets of $t$ elements that are distinct from each other). First, we note the following
@@
|$\ket{\mathcal{A}_{t}^{\mathsf{FHFHFO}\cdot G}}=\sqrt{\frac{(2^{n}-t)!}{2^{n}!}}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{t}\in(\{0,1\}^{n})^{t}\\ y_{1},\ldots,y_{t}\in(\{0,1\}^{n})^{t}\end{subarray}}\left(\prod_{i=1}^{t}\left(\ket{y_{i}}\!\!\bra{x_{i}}\cdot G\cdot A_{i}\right)\ket{0}\right)\otimes\ket{\mathsf{fhfhf}_{\{(x_{i},y_{i})\}_{i=1}^{t}}}\,.$
@@
$\mathsf{Biject}\cdot\ket{\mathcal{A}_{t}^{\mathsf{FHFHFO}\cdot G}}=\sqrt{\frac{(2^{n}-t)!}{2^{n}!}}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{t}\in(\{0,1\}^{n})^{t}_{\mathrm{dist}}\\ y_{1},\ldots,y_{t}\in(\{0,1\}^{n})^{t}_{\mathrm{dist}}\end{subarray}}\left(\prod_{i=1}^{t}\left(\ket{y_{i}}\!\!\bra{x_{i}}\cdot G\cdot A_{i}\right)\ket{0}\right)\otimes\ket{\mathsf{fhfhf}_{\{(x_{i},y_{i})\}_{i=1}^{t}}}\,.$
It is clear that applying $\mathsf{Compress}$ on this state yields the same state as $\Pi^{(t)}\ket{\mathcal{A}_{t}^{\mathsf{PR}\cdot G}}$, as the $\mathsf{FHF}$-relation states become their actual relations, and the binary phase gets canceled by the phase of the $\mathsf{Compress}$ oracle. ∎
@@
$\mathrm{td}\left(\mathbb{E}_{\mathsf{C}}|\mathcal{A}^{\mathsf{FHFHFO}\cdot\mathsf{C}}\rangle\!\langle\mathcal{A}^{\mathsf{FHFHFO}\cdot\mathsf{C}}|,\mathsf{Bijection}\cdot\mathbb{E}_{\mathsf{C}}|\mathcal{A}^{\mathsf{FHFHFO}\cdot\mathsf{C}}\rangle\!\langle\mathcal{A}^{\mathsf{FHFHFO}\cdot\mathsf{C}}|\cdot\mathsf{Bijection}\right)\leq O\left(\frac{t^{2}}{2^{n}}\right)\,.$
@@
Following the proof in [ma2024pseudorandom], we observe that $\mathsf{Compress}^{\dagger}\cdot\mathsf{Compress}\cdot\mathsf{Bijection}=\mathsf{Bijection}$. Then applying Lemma 3.2, we have the following bound on the trace distance:
$1-\mathrm{Tr}\left(\mathsf{Bijection}\cdot\mathbb{E}_{\mathsf{C}}|\mathcal{A}^{\mathsf{FHFO}\cdot\mathsf{C}}\rangle\!\langle\mathcal{A}^{\mathsf{FHFO}\cdot\mathsf{C}}|\cdot\mathsf{Bijection}\right)$
$\hskip 42.67912pt=1-\mathrm{Tr}\left(\mathsf{Compress}\cdot\mathsf{Bijection}\cdot\mathbb{E}_{\mathsf{C}}|\mathcal{A}^{\mathsf{FHFHFO}\cdot\mathsf{C}}\rangle\!\langle\mathcal{A}^{\mathsf{FHFHFO}\cdot\mathsf{C}}|\cdot\mathsf{Bijection}\cdot\mathsf{Compress}^{\dagger}\right)$
$\hskip 42.67912pt=1-\mathrm{Tr}\left(\Pi^{(t)}\cdot\mathbb{E}_{\mathsf{C}}|\mathcal{A}^{\mathsf{PR}\cdot\mathsf{C}}\rangle\!\langle\mathcal{A}^{\mathsf{PR}\cdot\mathsf{C}}|\cdot\Pi^{(t)}\right)$
$\hskip 42.67912pt\leq\frac{t(t-1)}{2^{n}}\,.$
This completes the proof of the lemma. ∎
Now, we put everything together, using the fact that [ma2024pseudorandom] has already shown that the path recording oracle itself is both right-invariant and indistinguishable from a Haar random unitary.
Theorem 6.25.
Let $n\in\mathbb{N}$ be the number of qubits as input to a quantum adversary. Then for all $t$-query adversaries $\mathcal{A}^{(\cdot)}$ making queries to a $n$ qubit oracle, the following holds
$\mathrm{td}\left(\mathbb{E}_{U\sim\mu(2^{n})}\left[\mathcal{A}^{U}(\ket{0})\right],\mathbb{E}_{\mathcal{O}\sim\mathsf{FHFHFC}}\left[\mathcal{A}^{\mathcal{O}}(\ket{0})\right]\right)\leq O\left(\frac{t^{2}}{2^{n}}\right)\,.$
Proof.
@@
    The adversary’s view of purification of the $\mathsf{FHFHFC}$ ensemble, after projecting onto distinct relation states.
@@
    The adversary’s view of the purification of the $\mathsf{PR}\cdot\mathsf{C}$ oracle, after projection onto bijective relations.
@@ 6.6.3 Instantiation under $\mathsf{HPS}$
Let $\vec{A}=\{\vec{A}_{\ell}\sim\mathbb{Z}_{2}^{m\times n}\}_{\ell=1}^{6}$, and let $\vec{k}=\{\vec{k}_{\ell}=(\theta_{1},\ldots,\theta_{m})\}_{\ell=1}^{5}$ be chosen uniformly at random (according to some discretization), with $m\geq n+\omega(\log n)$. We conjecture that the following gives rise to a pseudorandom unitary:
@@
Of course, we note that this conjecture is an assumption about the distinguishing advantage of any computationally bounded adversary. Our evidence that this conjecture is true comes from the security of the repeated DH pseudo-random unitary that we proved in the previous section. In that section, we took “ideal” HPS instances for all the diagonal gates, which can be viewed as running the HPS time evolution for an exponential time. Thus, our conjecture essentially ammounts to the fact that stopping this evolution by the HPS unitary after some polynomial-amount of time yielsd a pseudorandom unitary instead of a truly Haar random one.

### v2

@@ 6.6.1 Security of the $\mathsf{FHFHFC}$ pseudo-random unitary
$\frac{1}{2^{n/2}}\sum_{w,y\in\{0,1\}^{n}}(-1)^{\braket{y,x\oplus w}}\ket{y}\frac{1}{2^{3n}}\sum_{\begin{subarray}{c}\theta_{1}^{(1)},\ldots,\theta_{2^{n}}^{(1)}\\ \theta_{1}^{(2)},\ldots,\theta_{2^{n}}^{(2)}\\ \theta_{1}^{(3)},\ldots,\theta_{2^{n}}^{(3)}\end{subarray}}\alpha_{R}^{\theta^{(1)},\theta^{(2)},\theta^{(3)}}\cdot\\ \exp\left(i\sum_{z}\theta_{z}^{(1)}(-1)^{\braket{x,z}}+\theta_{z}^{(2)}(-1)^{\braket{w,z}}+\theta_{z}^{(3)}(-1)^{\braket{y,z}}\right)\ket{\theta^{(1)},\theta^{(2)},\theta^{(3)}}\,.$
Expanding out the definition of $\alpha_{R}^{\theta^{(1)},\theta^{(2)},\theta^{(3)}}$, pushing the sum over $w$ into the purifying register, and combining the product of exponentials yields the desired state. ∎
A subtle difference between this version and Claim $5$ of [ma2024pseudorandom] is the difference in normalization. However, since the size of the relation is polynomial in $n$, this will contribute a trace distance error of at most $t^{2}/2^{n}$, which we account for later.
Now, we have shown that the relation states for the $\mathsf{FHFHF}$ oracle are orthogonal whenever the relation is collision free, and that the action of the $\mathsf{FHF}$ oracle is to output a random $y$ and store $(x,y)$ in the relation state. Now, we make the connection to the path recording oracle formal by defining two isometries that will act on the purifying register, $\mathsf{Compress}$ and $\mathsf{Collide}$. The compress isometry acts as follows:
$\mathsf{Compress}=\sum_{R\ \mathrm{collision\ free}}\ket{R}\!\!\bra{\mathsf{fhfhf}_{R}}\,.$ | .
@@
$\mathsf{Collide}=\sum_{R\ \mathrm{collision\ free}}|\mathsf{fhfhf}_{R}\rangle\!\langle\mathsf{fhfhf}_{R}|\,.$
Let $\Pi^{(t)}$ be the projection onto bijection relation states, as defined in [ma2024pseudorandom], and $\Lambda^{(t)}$ be the projection onto collision free relation states. Then we have the following lemma.
@@
$\mathrm{td}\left(\mathsf{Compress}\cdot\mathsf{Collide}\cdot|\mathcal{A}_{t}^{\mathsf{FHFHFO}\cdot G}\rangle\!\langle\mathcal{A}_{t}^{\mathsf{FHFHFO}\cdot G}|\cdot\mathsf{Collide}\cdot\mathsf{Compress},\Lambda^{(t)}\cdot|\mathcal{A}^{\mathsf{PR}\cdot G}_{t}\rangle\!\langle\mathcal{A}^{\mathsf{PR}\cdot G}_{t}|\cdot\Lambda^{(t)}\right)\\ \leq O\left(\frac{t^{2}}{2^{n}}\right)\,.$
@@
First, note that collision-free relations are a subset of bijective relations, and the projectors $\Pi^{(t)}$ and $\Lambda^{(t)}$ commute, which $\Lambda^{(t)}=\Lambda^{(t)}\Pi^{(t)}$. Similar to the proof from [ma2024pseudorandom], we can explicitly write out the states |$\ket{\mathcal{A}_{t}^{\mathcal{O}}}$ in either case. For a set $S$, let $S^{t}$ denote the $t$-fold cartestian product of the set, and $S^{t}_{\mathrm{dist}}$ be the distinct subset of $S^{t}$ (i.e. all sets of $t$ elements that are distinct from each other). First, we note the following
@@
Further performing the projection onto collision free relations yields the following:
$\Pi^{(t)}\cdot\ket{\mathcal{A}_{t}^{\mathsf{PR}\cdot G}}=\sqrt{\frac{(2^{n}-t)!}{2^{n}!}}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{t}\in(\{0,1\}^{n})^{t}_{\mathrm{dist}}\\ y_{1},\ldots,y_{t}\in\mathsf{CF}_{x_{1},\ldots,x_{t}}\end{subarray}}\left(\prod_{i=1}^{t}\left(\ket{y_{i}}\!\!\bra{x_{i}}\cdot G\cdot A_{i}\right)\ket{0}\right)\otimes\ket{\{(x_{i},y_{i})\}_{i=1}^{t}}\,.$
@@
|$\ket{\mathcal{A}_{t}^{\mathsf{FHFHFO}\cdot G}}=\sqrt{\frac{1}{2^{nt}}}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{t}\in(\{0,1\}^{n})^{t}\\ y_{1},\ldots,y_{t}\in(\{0,1\}^{n})^{t}\end{subarray}}\left(\prod_{i=1}^{t}\left(\ket{y_{i}}\!\!\bra{x_{i}}\cdot G\cdot A_{i}\right)\ket{0}\right)\otimes\ket{\mathsf{fhfhf}_{\{(x_{i},y_{i})\}_{i=1}^{t}}}\,.$
@@
$\mathsf{Collide}\cdot\ket{\mathcal{A}_{t}^{\mathsf{FHFHFO}\cdot G}}=\sqrt{\frac{1}{2^{nt}}}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{t}\in(\{0,1\}^{n})^{t}_{\mathrm{dist}}\\ y_{1},\ldots,y_{t}\in\mathsf{CF}_{x_{1},\ldots,x_{t}}\end{subarray}}\left(\prod_{i=1}^{t}\left(\ket{y_{i}}\!\!\bra{x_{i}}\cdot G\cdot A_{i}\right)\ket{0}\right)\otimes\ket{\mathsf{fhfhf}_{\{(x_{i},y_{i})\}_{i=1}^{t}}}\\ +\mathsf{Collide}\cdot\sqrt{\frac{1}{2^{nt}}}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{t}\in(\{0,1\}^{n})^{t}_{\mathrm{dist}}\\ y_{1},\ldots,y_{t}\not\in\mathsf{CF}_{x_{1},\ldots,x_{t}}\end{subarray}}\left(\prod_{i=1}^{t}\left(\ket{y_{i}}\!\!\bra{x_{i}}\cdot G\cdot A_{i}\right)\ket{0}\right)\otimes\ket{\mathsf{fhfhf}_{\{(x_{i},y_{i})\}_{i=1}^{t}}}\,.$
Since we showed that collision free relation states are orthogonal to each other, compress will map the first term to the correct relation states, and shrink the trace of the second term. However, since the probability that a $y$ is not collision free relative to $x$ is at most $t^{3}/2^{n}$, the trace of the second term is at most $t^{3}/2^{n}$, and the normalization difference in the first term contributes at most a $t^{2}/2^{n}$ trace distance error. Applying the triangle inequality on these two steps, we get the desired error bound. ∎
Next, we show that the states after projecting onto $\Pi^{(t)}$ (bijections) and $\Lambda^{(t)}$ (collision free relations) are close to each other in trace distance.
Lemma 6.32.
For all $t$-query adversaries $\mathcal{A}_{t}^{\mathcal{O}}$ and unitaries $G$, the following bound holds:
$\mathrm{td}\left(\Lambda^{(t)}\cdot|\mathcal{A}^{\mathsf{PR}\cdot G}\rangle\!\langle\mathcal{A}^{\mathsf{PR}\cdot G}|\cdot\Lambda^{(t)},\Pi^{(t)}\cdot|\mathcal{A}^{\mathsf{PR}\cdot G}\rangle\!\langle\mathcal{A}^{\mathsf{PR}\cdot G}|\cdot\Pi^{(t)}\right)\leq O\left(\sqrt{\frac{t^{3}}{2^{n}}}\right)\,.$
Proof.
Since the probability that a uniformly random $y_{1},\ldots,y_{t}$ is collision free is at least $1-t^{3}/2^{n}$ when measuring a uniformly random distinct $y_{1},\ldots,y_{t}$, applying the gentle measurement measurement lemma yields the desired trace distance bound. ∎
@@
$\mathrm{td}\left(\mathbb{E}_{\mathsf{C}}\left[|\mathcal{A}^{\mathsf{FHFHFO}\cdot\mathsf{C}}\rangle\!\langle\mathcal{A}^{\mathsf{FHFHFO}\cdot\mathsf{C}}|\right],\mathsf{Collide}\cdot\mathbb{E}_{\mathsf{C}}\left[|\mathcal{A}^{\mathsf{FHFHFO}\cdot\mathsf{C}}\rangle\!\langle\mathcal{A}^{\mathsf{FHFHFO}\cdot\mathsf{C}}|\right]\cdot\mathsf{Collide}\right)\\ \leq O\left(\sqrt{\frac{t^{3}}{2^{n}}}\right)\,.$
@@
Following the proof in [ma2024pseudorandom], we observe that $\mathsf{Compress}^{\dagger}\cdot\mathsf{Compress}\cdot\mathsf{Collide}=\mathsf{Collide}$. Then applying Lemma 3.2, we have the following bound on the trace distance:
$1-\mathrm{Tr}\left(\mathsf{Collide}\cdot\mathbb{E}_{\mathsf{C}}\left[|\mathcal{A}^{\mathsf{FHFO}\cdot\mathsf{C}}\rangle\!\langle\mathcal{A}^{\mathsf{FHFO}\cdot\mathsf{C}}|\right]\cdot\mathsf{Collide}\right)$
$\hskip 42.67912pt=1-\mathrm{Tr}\left(\mathsf{Compress}\cdot\mathsf{Collide}\cdot\mathbb{E}_{\mathsf{C}}\left[|\mathcal{A}^{\mathsf{FHFHFO}\cdot\mathsf{C}}\rangle\!\langle\mathcal{A}^{\mathsf{FHFHFO}\cdot\mathsf{C}}|\right]\cdot\mathsf{Collide}\cdot\mathsf{Compress}^{\dagger}\right)$
$\hskip 42.67912pt\leq 1-\mathrm{Tr}\left(\Lambda^{(t)}\cdot\mathbb{E}_{\mathsf{C}}\left[|\mathcal{A}^{\mathsf{PR}\cdot\mathsf{C}}\rangle\!\langle\mathcal{A}^{\mathsf{PR}\cdot\mathsf{C}}|\right]\cdot\Lambda^{(t)}\right)+O\left(\sqrt{\frac{t^{3}}{2^{n}}}\right)$
$\hskip 42.67912pt\leq 1-\mathrm{Tr}\left(\Pi^{(t)}\cdot\mathbb{E}_{\mathsf{C}}\left[|\mathcal{A}^{\mathsf{PR}\cdot\mathsf{C}}\rangle\!\langle\mathcal{A}^{\mathsf{PR}\cdot\mathsf{C}}|\right]\cdot\Pi^{(t)}\right)+O\left(\sqrt{\frac{t^{3}}{2^{n}}}\right)$
$\hskip 42.67912pt\leq\frac{t(t-1)}{2^{n}}+O\left(\sqrt{\frac{t^{3}}{2^{n}}}\right)\,.$
Here the second line uses Lemma 6.31, the second to last line uses the previous lemma, which states that the projection onto collision free relations is close to the projection onto bijections, and the final line uses Corollary 4.1 from [ma2024pseudorandom]. This completes the proof of the lemma. ∎
Now, we put everything together and prove Theorem 6.17, using the fact that [ma2024pseudorandom] has already shown that the path recording oracle itself is both right-invariant and indistinguishable from a Haar random unitary.
Proof of Theorem 6.17.
@@
    The adversary’s view of purification of the $\mathsf{FHFHFC}$ ensemble, after projecting onto collision free relation states.
@@
    The adversary’s view of the purification of the $\mathsf{PR}\cdot\mathsf{C}$ oracle, after projection onto collision free relations.
@@
Remark 6.34.
We note that the proof in this section also applies to the case when $\mathsf{F}$ is a binary phase state. The only property we used was the orthogonality of set states, which also applies to the set states formed using the purification technique to binary phase states.
Conjecture 6.35.
We conjecture that it can be shown that $\mathsf{FHFHF}$ itself (without the random Clifford operation) for uniformly random phases is a strong (i.e. inverse secure) pseudorandom unitary as well. As noted before, proving this likely requires understanding when $((x_{i},y_{i}))_{i}$ is uniquely determined by the set of $x$’s, $y$’s, and $x\oplus y$’s.
@@ 6.6.3 Instantiation under $\mathsf{HPS}$
Let $\vec{A}=\{\vec{A}_{\ell}\sim\mathbb{Z}_{2}^{m\times n}\}_{\ell=1}^{6}$, and let $\vec{k}=\{\vec{k}_{\ell}=(\theta_{1},\ldots,\theta_{m})\}_{\ell=1}^{6}$ be chosen uniformly at random (according to some discretization), with $m\geq n+\log(n)$. We conjecture that the following gives rise to a pseudorandom unitary:
@@
Of course, we note that this conjecture is an assumption about the distinguishing advantage of any computationally bounded adversary. Our evidence that this conjecture is true comes from the security of the repeated DH pseudo-random unitary that we proved in the previous section. In that section, we took “ideal” HPS instances (i.e. completely random diagonal unitary) for all the diagonal gates, which can be viewed as running an exponentially deep HPS circuit. Thus, our conjecture essentially amounts to the fact that stopping this evolution by the HPS unitary after some polynomial-amount of time yields a pseudorandom unitary instead of a truly Haar random one.